# Beyond Statistical Estimation: Differentially Private Individual Computation via Shuffling

### v2

@@ V-C Discussion on Post-computation Communication
The post-computation communication may also have privacy implications for other users not in the matched pair. Privacy amplification via shuffling against an adversary relies on the number of parties that remain anonymous. Recall that in Equation 1, the amplified $\epsilon_{c}$ relies on the number $n^{\prime}_{i^{*}}=|G_{i^{*}}-C_{i^{*}}|$ of uncorrupted users in a particular group $G_{i^{*}}$. From the perspective of the adversary, if the post-computation communication compromises the anonymity of an additional set $U$ of the users in $G_{i^{*}}$, then $n^{\prime}_{i^{*}}$ becomes $|G_{i^{*}}-C_{i^{*}}-U|$. Accordingly, the privacy amplification effect for users in $G_{i^{*}}$ that are still anonymous is weakened.
@@
VI Optimal Randomizers
VI-A Inadequacy of Existing Randomizers
@@
where $\mathcal{R}$ is the LDP-randomizer employed by the users. Additionally, in the PIC model, user data is typically multi-dimensional (e.g., location, gradient), making sanitization significantly more difficult compared to scalar data. Another notable characteristic of the PIC model is that the local differential privacy budget $\epsilon$ is relatively large, often scaling linearly with $\log(n^{\prime}_{i^{*}})$, as implied by Theorem V.2. These factors together create issues when existing LDP randomizers are applied directly in the PIC model.
@@
There have been LDP randomizers that submit complete reports. One obvious strategy is to explicitly split the local budget into $d$ parts and then apply a one-dimensional LDP mechanism independently to each dimension, or implicitly distribute the budget among dimensions, as seen in the Laplace [30], PlanarLaplace [2], PrivUnit [12], and PrivUnitG mechanisms [6]. However, these approaches are sub-optimal in the high budget regime. Specifically, even if we use the optimal one-dimension randomizer [7], splitting the budget across each dimension and then applying any randomizer for each dimension will result in a mean squared error (MSE) of at least $\frac{d}{(e^{\epsilon/d}-1)^{2/3}}$. The Laplace/PlanarLaplace mechanisms introduce an MSE rate of $\frac{d}{\epsilon^{2}}$, while the PrivUnit/PrivUnitG mechanisms incur an MSE rate of $\frac{d}{\min\{\epsilon,\epsilon^{2}\}}$. In contrast, later we will show that the MSE rate can be improved to $(e^{\epsilon}-1)^{-2/(d+2)}$ (see Section VI-C and Appendix -D).
@@
VI-B Randomizer Design
We now introduce an asymptotically optimal randomizer, tailored for the PIC model. The randomizer uses an LDP mechanism, which we termed as Minkowski Response. For ease of presentation, here we will focus primarily on the $\ell_{2}$ case (and the $\ell_{+\infty}$ case in Appendix -D), but the mechanism can be generalized to other Minkowski distances.
@@ VI-B Randomizer Design
For any input $x\in\mathbb{X}$, Minkowski Response outputs an output $y\in\mathbb{Y}_{r}$ with relatively high probability in the cap area $\mathbb{B}_{r}(x)$ and relatively low probability in remaining domain $\mathbb{Y}_{r}\backslash\mathbb{B}_{r}(x)$ (see Figure 4). Formally:
@@
It is obvious that the output of Minkowski Response is informative in every dimension. Therefore it avoids problems brought up by incomplete reports. Also, intuitively it has a better utility because the mechanism is more likely to output a value near the true value (i.e., in $\mathbb{B}_{r}(x)$), than from other parts of the output domain (i.e. $\mathbb{Y}_{r}\backslash\mathbb{B}_{r}(x)$). When the budget $\epsilon$ gets large, the error rate of Minkowski Response decays faster than in previous LDP mechanisms. More specifically, the decay rate of Minkowski Response is $(e^{\epsilon}-1)^{-2/(d+2)}$ (Equation 6 in Appendix -D), while that of the previous mechanisms is $d/(e^{\epsilon/d}-1)^{2/3}$ or $d/\epsilon^{2}$. Therefore, the utility advantage of Minkowski Response becomes more significant when $\epsilon$ gets larger. When $n^{\prime}_{i^{*}}\to+\infty$ (and thus $\epsilon\to+\infty$), $r$ becomes 0, and the error goes to zero (i.e. no additional error).
VI-C Analysis of Minkowski Response
The local privacy guarantee of the randomizer is presented in Theorem VI.1.
Theorem VI.1 (Local Privacy Guarantee)
@@ VI-C Analysis of Minkowski Response
Next, we analyze the utility of the Minkowski Response in the PIC model. As previously mentioned, in the PIC model, the single report error is a more appropriate measure of utility compared to the statistical errors used in the conventional shuffle model. In Theorem VI.2, we examine the single report error in the PIC model and establish its lower bound. The proof is provided in Appendix -C.
Theorem VI.2 (Error Lower Bounds)
Given $d\in\mathbb{N}$, $\epsilon_{c}>0$, $\delta\in(0,0.5]$, $\mathbb{X}=\mathbb{B}_{1}(\{0\}^{d})$, then for any randomizer $\mathcal{R}:\mathbb{X}\mapsto\mathbb{R}^{d^{\prime}}$ such that $\mathcal{S}\circ\mathcal{R}(X)$ and $\mathcal{S}\circ\mathcal{R}(X^{\prime})$ are $(\epsilon_{c},\delta)$-indistinguishable for all possible neighboring datasets $X,X^{\prime}\in\mathbb{X}^{n}$, and for any estimator $f:\mathbb{R}^{d^{\prime}}\mapsto\mathbb{R}^{d}$, we have:
@@
The above theorem suggests that in the PIC model, for any randomizer, the single report error is at least $\tilde{\Omega}\big{(}{1}/{n^{\frac{2}{d+2}}}\big{)}$. Clearly, a randomizer offers better utility if its error is closer to this bound. For the randomizer using Minkowski Response, as presented in section VI-B, we can demonstrate that its single report error has an upper bound. This is summarized in Theorem VI.3, with the proof provided in Appendix -D.
Theorem VI.3 (Error Upper Bounds)
@@
We observe that in most applications, $(\epsilon_{c},\delta)$ are given as fixed system parameters. If we consider $(\epsilon_{c},\delta)$ as constants, then the error upper bound of the Minkowski Response randomizer in Theorem VI.3 is $\tilde{O}\left(\frac{1}{n^{2/(d+2)}}\right)$, which matches the error lower bound of the PIC model in Theorem VI.2. This implies that the utility of the Minkowski Response is asymptotically optimal. In contrast, using existing LDP randomizers in the PIC model results in a larger single report error of $\tilde{O}\left(\frac{d}{n^{2/(3d)}}\right)$ or $\tilde{O}\left(\frac{d}{\log^{2}n}\right)$ (note that $d>1$ and $n$ is often not small). Although asymptotic notations describe behavior as $n\rightarrow\infty$, in practice, the utility advantage of the Minkowski Response becomes noticeable without $n$ being very large: in our experiments, the Minkowski Response randomizer outperforms existing LDP randomizers in the PIC model once $n$ reaches the order of $10^{2}$. If we compare Minkowski Response in the PIC model to using LDP directly (without shuffling), the utility advantage is even greater: any randomizers in the LDP model must endure a single report error of $\Omega\left(\frac{d}{\epsilon_{c}^{2}}\right)$ when $\epsilon_{c}\leq O(1)$ [12, 28].
VII Experimental Evaluation
@@
VII-A Spatial Crowdsourcing
Datasets We use two real-world datasets: GMission dataset [18] for scientific simulation, and EverySender dataset [82] for campus-based micro-task completion. Details about the two datasets are summarized in Table II, including the number of users/workers, location domain range, and serving radius of workers about these two datasets (more information can be found in Appendix -F). We normalize location data to the domain $[-1,1]\times[-1,1]$ and scale the serving radius to $0.41.0\cdot\frac{1-(-1)}{5-0}=$ 0.4 correspondingly.
TABLE II: Summary statistics of spatial crowdsourcing datasets.
@@ VII-A Spatial Crowdsourcing
LDP randomizers We use the Minkowski response for location randomization. As a comparison, in the local model of DP (e.g., in [85, 80, 84]), we compare with existing mechanisms including Laplace [31], Staircase [38], PlanarLaplace with geo-indistinguishability [2], SquareWave [58], PrivUnit [12] and its Gaussian variant PrivUnitG in [6].
@@
Finally, we present the utility comparison using the maximum matching algorithm, as shown in Fig. 7. In this case, task-specific utility is assessed by the successful matching ratio:
@@
VII-B Location-based Social Systems
TABLE III: Details of location social network datasets.
@@ VII-B Location-based Social Systems
Datasets We use two real-world datasets: Gowalla dataset [21] and Foursquare dataset [89]. Gowalla and Foursquare are location-based social network websites where users share their locations by checking-in. Details about the two datasets are summarized in Table III. As before, the location data is normalized to $[-1,1]\times[-1,1]$.
LDP randomizers The LDP randomizers used are the same as those in Section VII-A.
Server-side algorithm The server performs the radius-based nearest neighbor (NN) search for the users, which is a common task in location-based social networks [21]. We set the search radius $\tau$ to 0.2, so that each user has several hundreds or thousands of neighbors, varying due to check-in densities. Note that in this application scenario, the returned neighbors may be deanonymized in the post-computation phase. Hence, the actual privacy guarantee in the PIC model depends on the number of users who remain anonymous (see discussion in Section V-C). Considering this, and each user normally has no more than $n\cdot\frac{\pi\tau^{2}}{2^{2}}<n\cdot 3.2\%$ neighbors, we use privacy amplification population $\lfloor n\cdot 90\%\rfloor$ instead of the group size $n$ when calculating the local budget given the global $\epsilon_{c}$.
@@
The task-specific utility metric we use for nearest neighbor queries is the $F_{1}$ score. Let $N_{i}$ denote the set of true neighbors of user $i$ within an $\ell_{2}$-distance $\tau$, and let $\widehat{N}_{i}$ denote the retrieved neighbor set computed using the sanitized reports. The precision of nearest neighbor queries is defined as $\frac{\sum_{i\in[n]}|N_{i}\cap\widehat{N}_{i}|}{\sum_{i\in[n]}|\widehat{N}_{i}|}$, and the recall is defined as $\frac{\sum_{i\in[n]}|N_{i}\cap\widehat{N}_{i}|}{\sum_{i\in[n]}|N_{i}|}$. The $F_{1}$ score is then calculated as:
@@
VII-C Federated Learning with Incentives
@@
Datasets We utilize the MNIST handwritten digit dataset to train a simple neural network using federated learning. The MNIST dataset comprises 60,000 images, with 50,000 designated as training samples. Each user receives one training sample and trains a neural network model, as detailed in Table IV, which contains $d=4292$ trainable parameters. In each round, $s=10,000$ users are randomly selected. These selected users locally compute the gradient vector using Stochastic Gradient Descent (SGD), then subsample 0.15% of the gradient vector dimensions (with unsampled dimensions set to zero), and clip the gradient values at a threshold of $c=0.00015$. Each user submits a sanitized version of the subsampled, clipped gradient vector as their report.
TABLE IV: Model architecture of the neural network.
@@
Server-side algorithm The server performs the following steps: first, it takes a sanitized gradient vector and uses randomizer-specific algorithms to estimate the true values of the sampled dimensions. These vectors with estimated values are denoted as $\hat{g}_{i}$. The server then aggregates the $\hat{g}_{i}$ vectors into a global gradient vector using a simple sum-and-average method. This global gradient vector is published so that users can update their local models. Additionally, the server computes the Shapley value for each user, which measures the marginal contribution of each $\hat{g}_{i}$ (see Appendix -H for details).
Experimental results We train each model for 80 rounds using Federated Learning. The utility comparison results are summarized in Tables V and VI, each reflecting a different global privacy budget. In these tables, we present the single report utility (Gradient $\ell_{2}$ error) and the task-specific utility (Shapley $\ell_{2}$ error). The Gradient $\ell_{2}$ error is calculated using the estimated gradients $\hat{g}_{i}$ and the true gradients $g_{i}$ (clipped and with all unsampled dimensions set to 0): $\frac{\sum_{i\in S}\|\hat{g}_{i}-{g}_{i}\|_{2}}{|S|}$, where $S$ is the set of randomly selected users. The Shapley $\ell_{2}$ error is computed as follows:$\frac{\sum\nolimits_{i\in S}\|\widehat{\texttt{Shapley}}_{i}-{\texttt{Shapley}}_{i}\|_{2}}{|S|}$, where $\widehat{\texttt{Shapley}}_{i}$ is derived from $\hat{g}_{i}$ and ${\texttt{Shapley}}_{i}$ is obtained from the unsanitized $g_{i}$. For reference, we also include the final models’ accuracy in the tables.
TABLE V: Utility comparison of federated learning with incentives: global privacy budget $(1,0.01/50000)$.
@@
TABLE VI: Utility comparison of federated learning with incentives: global privacy budget $(3,0.01/50000)$.
@@
We observe in the tables that the utility in the LDP model is significantly worse than in the PIC model, as expected. This is consistent across all metrics and the final model accuracy. In the local model, the performance of Staircase, Squarewave, and Minkowski mechanisms is poorer compared to the others, but for different reasons. Staircase and Squarewave are designed for single-dimensional data, requiring the privacy budget to be split across dimensions when sanitizing vectors, which leads to reduced utility. For Minkowski, the issue lies in the 80-round federated learning process, where the privacy budget for each round is relatively small (e.g., between 0.2 and 0.75). As previously mentioned, the Minkowski mechanism is intended to achieve better utility with a large local privacy budget. In small privacy budget scenarios, it offers no advantage and may even perform worse. In the PIC model, the performance of Staircase and Squarewave remains poor. However, Minkowski now performs the best among all randomizers. This improvement is due to the privacy amplification effect in PIC, which increases the local privacy budget for each round to between 4.0 and 6.0.
VIII Conclusion

### v4

@@ 5.3 Discussion on Post-computation Communication
[…]
@@
6 Optimal Randomizers
6.1 Inadequacy of Existing Randomizers
@@
[…]
@@
There have been LDP randomizers that submit complete reports. One obvious strategy is to explicitly split the local budget into $d$ parts and then apply a one-dimensional LDP mechanism independently to each dimension, or implicitly distribute the budget among dimensions, as seen in the Laplace [30], PlanarLaplace [2], PrivUnit [12], and PrivUnitG mechanisms [6]. However, these approaches are sub-optimal in the high budget regime. Specifically, even if we use the optimal one-dimension randomizer [7], splitting the budget across each dimension and then applying any randomizer for each dimension will result in a mean squared error (MSE) of at least $\frac{d}{(e^{\epsilon/d}-1)^{2/3}}$. The Laplace/PlanarLaplace mechanisms introduce an MSE rate of $\frac{d}{\epsilon^{2}}$, while the PrivUnit/PrivUnitG mechanisms incur an MSE rate of $\frac{d}{\min\{\epsilon,\epsilon^{2}\}}$. In contrast, later we will show that the MSE rate can be improved to $(e^{\epsilon}-1)^{-2/(d+2)}$ (see Section 6.3 and Appendix A).
@@
6.2 Randomizer Design
We now introduce an asymptotically optimal randomizer, tailored for the PIC model. The randomizer uses an LDP mechanism, which we termed as Minkowski Response. For ease of presentation, here we will focus primarily on the $\ell_{2}$ case (and the $\ell_{+\infty}$ case in Appendix A), but the mechanism can be generalized to other Minkowski distances.
@@ 6.2 Randomizer Design
[…]
@@
It is obvious that the output of Minkowski Response is informative in every dimension. Therefore it avoids problems brought up by incomplete reports. Also, intuitively it has a better utility because the mechanism is more likely to output a value near the true value (i.e., in $\mathbb{B}_{r}(x)$), than from other parts of the output domain (i.e. $\mathbb{Y}_{r}\backslash\mathbb{B}_{r}(x)$). When the budget $\epsilon$ gets large, the error rate of Minkowski Response decays faster than in previous LDP mechanisms. More specifically, the decay rate of Minkowski Response is $(e^{\epsilon}-1)^{-2/(d+2)}$ (Equation 4 in Appendix A), while that of the previous mechanisms is $d/(e^{\epsilon/d}-1)^{2/3}$ or $d/\epsilon^{2}$. Therefore, the utility advantage of Minkowski Response becomes more significant when $\epsilon$ gets larger. When $n^{\prime}_{i^{*}}\to+\infty$ (and thus $\epsilon\to+\infty$), $r$ becomes 0, and the error goes to zero (i.e. no additional error).
6.3 Analysis of Minkowski Response
The local privacy guarantee of the randomizer is presented in Theorem 6.1.
Theorem 6.1 (Local Privacy Guarantee).
@@ 6.3 Analysis of Minkowski Response
[…]
Theorem 6.2 (Error Lower Bounds).
[…]
@@
[…]
Theorem 6.3 (Error Upper Bounds).
@@
We observe that in most applications, $(\epsilon_{c},\delta)$ are given as fixed system parameters. If we consider $(\epsilon_{c},\delta)$ as constants, then the error upper bound of the Minkowski Response randomizer in Theorem 6.3 is $\tilde{O}\left(\frac{1}{n^{2/(d+2)}}\right)$, which matches the error lower bound of the PIC model in Theorem 6.2. This implies that the utility of the Minkowski Response is asymptotically optimal. In contrast, using existing LDP randomizers in the PIC model results in a larger single report error of $\tilde{O}\left(\frac{d}{n^{2/(3d)}}\right)$ or $\tilde{O}\left(\frac{d}{\log^{2}n}\right)$ (note that $d>1$ and $n$ is often not small). Although asymptotic notations describe behavior as $n\rightarrow\infty$, in practice, the utility advantage of the Minkowski Response becomes noticeable without $n$ being very large: in our experiments, the Minkowski Response randomizer outperforms existing LDP randomizers in the PIC model once $n$ reaches the order of $10^{2}$. If we compare Minkowski Response in the PIC model to using LDP directly (without shuffling), the utility advantage is even greater: any randomizers in the LDP model must endure a single report error of $\Omega\left(\frac{d}{\epsilon_{c}^{2}}\right)$ when $\epsilon_{c}\leq O(1)$ [12, 28].
7 Experimental Evaluation
@@
7.1 Spatial Crowdsourcing
Datasets We use two real-world datasets: GMission dataset [18] for scientific simulation, and EverySender dataset [86] for campus-based micro-task completion. Details about the two datasets are summarized in Table 2, including the number of users/workers, location domain range, and serving radius of workers about these two datasets (more information can be found in Appendix C of the full version [91]). We normalize location data to the domain $[-1,1]\times[-1,1]$ and scale the serving radius to $0.41.0\cdot\frac{1-(-1)}{5-0}=$ 0.4 correspondingly.
Table 2: Summary statistics of spatial crowdsourcing datasets.
@@ 7.1 Spatial Crowdsourcing
[…]
@@
Finally, we present the utility comparison using the maximum matching algorithm, as shown in Fig. 7. In this case, utility is assessed by the successful matching ratio:
@@
7.2 Location-based Social Systems
Table 3: Details of location social network datasets.
@@ 7.2 Location-based Social Systems
Datasets We use two real-world datasets: Gowalla dataset [21] and Foursquare dataset [94]. Gowalla and Foursquare are location-based social network websites where users share their locations by checking-in. Details about the two datasets are summarized in Table 3. As before, the location data is normalized to $[-1,1]\times[-1,1]$.
LDP randomizers The LDP randomizers used are the same as those in Section 7.1.
Server-side algorithm The server performs the radius-based nearest neighbor (NN) search for the users, which is a common task in location-based social networks [21]. We set the search radius $\tau$ to 0.2, so that each user has several hundreds or thousands of neighbors, varying due to check-in densities. Note that in this application scenario, the returned neighbors may be deanonymized in the post-computation phase. Hence, the actual privacy guarantee in the PIC model depends on the number of users who remain anonymous (see discussion in Section 5.3). Considering this, and each user normally has about $n\cdot\frac{\pi\cdot\tau^{2}}{2^{2}}\approx n\cdot 3.2\%$ neighbors, we use privacy amplification population $\lfloor n\cdot 90\%\rfloor$ instead of the group size $n$ when calculating the local budget given the global $\epsilon_{c}$.
@@
The task-specific utility metric we use for nearest neighbor queries is the $F_{1}$ score. Let $N_{i}$ denote the set of true neighbors of user $i$ within an $\ell_{2}$-distance $\tau$, and let $\widehat{N}_{i}$ denote the retrieved neighbor set computed using the sanitized reports. The precision of nearest neighbor queries is defined as ${\sum_{i\in[n]}|N_{i}\cap\widehat{N}_{i}|}/{\sum_{i\in[n]}|\widehat{N}_{i}|}$, and the recall is defined as ${\sum_{i\in[n]}|N_{i}\cap\widehat{N}_{i}|}/{\sum_{i\in[n]}|N_{i}|}$. The $F_{1}$ score is as:
@@
7.3 Federated Learning with Incentives
@@
Datasets We utilize the MNIST handwritten digit dataset to train a simple neural network using federated learning. The MNIST dataset comprises 60,000 images, with 50,000 designated as training samples. Each user receives one training sample and trains a neural network model, as detailed in Table 4, which contains $d=4292$ trainable parameters. In each round, $s=10,000$ users are randomly selected. These selected users locally compute the gradient vector using Stochastic Gradient Descent (SGD), then subsample 0.15% of the gradient vector dimensions (with unsampled dimensions set to zero), and clip the gradient values at a threshold of $c=0.00015$. Each user submits a sanitized version of the subsampled, clipped gradient vector as their report.
Table 4: Model architecture of the neural network.
@@
Server-side algorithm The server performs the following steps: first, it takes a sanitized gradient vector and uses randomizer-specific algorithms to estimate the true values of the sampled dimensions. These vectors with estimated values are denoted as $\hat{g}_{i}$. The server then aggregates the $\hat{g}_{i}$ vectors into a global gradient vector using a simple sum-and-average method. This global gradient vector is published so that users can update their local models. Additionally, the server computes the Shapley value for each user, which measures the marginal contribution of each $\hat{g}_{i}$ (see Appendix E for details in the full version [91]).
Experimental results We train each model for 80 rounds using Federated Learning. The utility comparison results are summarized in Tables 5 and 6, each reflecting a different global privacy budget. In these tables, we present the single report utility (Gradient $\ell_{2}$ error) and the task-specific utility (Shapley $\ell_{2}$ error). The Gradient $\ell_{2}$ error is calculated using the estimated gradients $\hat{g}_{i}$ and the true gradients $g_{i}$ (clipped and with all unsampled dimensions set to 0): $\frac{\sum_{i\in S}\|\hat{g}_{i}-{g}_{i}\|_{2}}{|S|}$, where $S$ is the set of randomly selected users. The Shapley $\ell_{2}$ error is computed as follows:$\frac{\sum\nolimits_{i\in S}\|\widehat{\texttt{Shapley}}_{i}-{\texttt{Shapley}}_{i}\|_{2}}{|S|}$, where $\widehat{\texttt{Shapley}}_{i}$ is derived from $\hat{g}_{i}$ and ${\texttt{Shapley}}_{i}$ is obtained from the unsanitized $g_{i}$. For reference, we also include the final models’ accuracy in the tables.
Table 5: Utility comparison of federated learning with incentives: global privacy budget $(1,0.01/50000)$.
@@
Table 6: Utility comparison of federated learning with incentives: global privacy budget $(3,0.01/50000)$.
@@
We observe in the tables that the utility in the LDP model is significantly worse than in the PIC model, as expected. This is consistent across all metrics and the final model accuracy. In the local model, the performance of Staircase, Squarewave, and Minkowski mechanisms is poorer compared to the others, but for different reasons. Staircase and Squarewave are designed for single-dimensional data, requiring the privacy budget to be split across dimensions when sanitizing vectors, which leads to reduced utility. For Minkowski, the issue lies in the 80-round federated learning process, where the privacy budget for each round is relatively small (e.g., between 0.2 and 0.75). As previously mentioned, the Minkowski is intended to achieve better utility with a large local privacy budget. In small budget scenarios, it offers no advantage and may even perform worse. In the PIC model, the performance of Staircase and Squarewave remains poor. However, Minkowski now performs the best among all randomizers. This improvement is due to the privacy amplification in PIC, which increases the local budget for each round to between 4.0 and 6.0.
8 Discussions
[…]
Decentralized download-then-compute settings. While our protocol initially allows a server to perform the computation $f$, it is also possible for users to first download $L$ and then compute $f$ locally. This download-then-compute paradigm can be seen as a special decentralized case of our PIC model (i.e. the server simply computes $f(x)=x$). It would also rely on the one-time random key to perform post-computation communication. The PIC protocol offers several advantages over the download-then-compute paradigm: (1) PIC accommodates more tasks, e.g. those dependent on private server information (validation dataset in federated learning with incentives). (2) Download-then-compute requires public $f$, while PIC allows proprietary server-side algorithms. (3) Better privacy. E.g. in taxi-hailing, download-then-compute allows each user to compute all matching pairs and location distribution, while PIC publishes encrypted matching results (without leaking unnecessary information).
[…]
Necessity of grouping users. While many applications involve a single type of participant (e.g., social systems and federated learning), other tasks require interaction between multiple participant types, where the server must use the type information of participants to execute the algorithm $f$ (e.g., in taxi-hailing services with types of drivers and users). From a privacy-preservation standpoint, it is also possible to apply varying levels of privacy protection across different groups.
[…]
Worst-case leakage and privacy under active attacks. We assume that an attacker can corrupt and control a subset of parties, but the data of the uncorrupted users remains $(\epsilon_{c},\delta)$-private against the attacker. The specific value of $\epsilon_{c}$ depends on the number of uncorrupted users within the same group (see Theorem 5.2, Eq. 1). In the worst-case scenario, where all but one user are corrupted, the privacy guarantee is reduced to $\epsilon$-LDP, as provided by the local randomizer, i.e., there is no privacy amplification through shuffling.
Multiple executions and reuse keys. When multiple/sequential PIC executions are needed (e.g., in federated learning), a fresh one-time user key for each execution is required, ensuring the adversary cannot correlate user activities. According to Eq. 1 and sequential composition of DP [31], the overall privacy consumption for $k$ executions is $\tilde{O}(\sqrt{k\cdot e^{\epsilon}/{n_{i}^{*}}})$. Otherwise, reusing keys in $k$ executions will consume more privacy (to a level of shuffling $n_{i}^{*}$ messages each is $k\cdot\epsilon$-LDP, i.e., $\tilde{O}(\sqrt{e^{k\cdot\epsilon}/{n_{i}^{*}}})$).
[…]
Multi-message settings. The PIC model allows each client to send one single message, to align with most non-private applications. However, restricting each client to sending one message in the shuffle model has intrinsic privacy amplification and utility limitations (as seen in statistical analyses within the shuffle model [40]). Extending PIC model to multi-message settings is an interesting future direction, though it is quite challenging (e.g., in taxi-hailing services, the messages from a driver must appear independent, then each message may match a different customer, leading to an oversale).
9 Conclusion